which of the following is the energy of a possible excited state of a He^+ ion?
(a) -1.09*10^-18 J
(b)-8.72*10^-19J
(c)-4.36*10^-19J
(d)-5.45*10^-19J

Grade:12

1 Answers

Arun
6 years ago

The different energy levels of an atom are given by the equation:

En = -13.6 Z2/n2 eV

Z is the atomic number. For hydrogen, Z = 1.

n = 1,2,3… and so on such that the ground state has n = 1 and the second energy level (first excited state) has n = 2.

So, for first excited state, n = 2

E = -13.6/22 = -13.6/4 = - 3.4 eV

E = -5.45 * 10^-19 joules